# On the geometric phase and its role in the design of elastic topological materials

### v1

@@ 6 Topological classification of elastic waveguides
The first perspective establishes a classification scheme of the waveguides into two types, $\mathcal{W}_{O}$ and $\mathcal{W}_{E}$. Let the number of times the loop corresponding to a waveguide encircles the origin be denoted $\nu$. Then, $\mathcal{W}_{O}$ ($\mathcal{W}_{E}$) is the class of the waveguide for which $\nu$ is odd (even). Waveguides in class $\mathcal{W}_{O}$ ($\mathcal{W}_{E}$) acquire (do not acquire) a geometric phase of $\pi$. For example, in Fig. 8a, the waveguide corresponding to loop $L_{1}$ ($L_{2}$) is in the class $\mathcal{W}_{O}$ ($\mathcal{W}_{E}$), which (does not) acquires a geometric phase of $\pi$.
This classification scheme satisfies the physical requirements outlined in the previous section: (i) Geometric perturbations to a given waveguide manifest as changes to the corresponding loop. If these perturbations are small enough, such as perturbing loop $L_{1}$ to $L_{1}^{\prime}$ in Fig. 8a, the waveguide class does not change. (ii) If a perturbation does deform the waveguide into a different class, there is an intermediate waveguide whose loop passes through the origin. In other words, a perturbation deforming a waveguide into a different class must close the gap in the $\omega$-$z$ plot. For example, in Fig. 8a, perturbing the waveguide corresponding to loop $L_{1}$ into a waveguide corresponding to loop $L_{3}$ passes through an intermediate waveguide corresponding to loop $L_{2}$, for which the gap is closed in the $\omega$-$z$ plot. The physical significance of this transition is shown by the geometric phase acquired by a flexural wave propagating down the waveguide. Perturbing a waveguide into a different class changes the acquired geometric phase from 00 to $\pi$ or vice versa.
Further, this classification scheme is also topological in nature. Recall from Sec. 5.1 that waveguides in class $\mathcal{W}_{O}$ ($\mathcal{W}_{E}$) give rise to a nontrivial (trivial) fiber bundle resembling the edge of a Möbius strip shown in Fig. 7e (edges of a cylinder, Fig. 7f). Thus, the classification scheme essentially distinguishes between these trivial and nontrivial fiber bundles. As discussed using Figs. 7e,f, the topological invariant is the holonomy of the parallel transport along the base manifold, which is precisely the topological geometric phase.
@@ 6.2 Topological classification of elastic metamaterials
The above classification scheme is readily adapted to ETMs once the fiber bundle description is identified. 1D ETMs that are infinite and periodic along the $z$ direction are considered. Further, the unit cell of length $a$ is assumed to be symmetric about the $xz$ and $yz$ planes to prevent coupling between the flexural, torsional, and longitudinal modes. Such metamaterials admit modulated plane wave solutions [54, 66] $\mathbf{u}(x,y,z,t)=\mathbf{U}(x,y,z;k)e^{\mathrm{i}(kz-\omega t)}$ to the NL equations, where $\mathbf{U}(x,y,z;k)$ is a periodic function of $z$ called the Bloch wave mode. The Bloch wave mode $\mathbf{U}(x,y,z;k)$ and angular frequency $\omega(k)$ are computed by solving a wavenumber dependent eigenvalue problem. These properties of the ETM are graphically described by plotting the dispersion relation, which shows the angular frequency as a function of the wavenumber, e.g., Figs. 9d-f. Due to the spatial periodicity of the metamaterial, the solutions are periodic in $k$ with periodicity $2\pi/a$, i.e., $\mathbf{U}(x,y,z;k)=\mathbf{U}(x,y,z;k+2\pi/a)$ and $\omega(k)=\omega(k+2\pi/a)$. In other words, the wavenumber is a circle $S^{1}$ represented as the interval $[-\pi/a,\pi/a]$ with the endpoints glued together. The wavenumber defines a circular base manifold. At each wavenumber $k$, the Bloch wave mode $\mathbf{U}(x,y,z;k)$ has a phase degree of freedom, as it can be arbitrarily multiplied by a phase factor $e^{\mathrm{i}\theta}$. This is described as a $U(1)$ fiber. Thus, a Bloch wave mode of an ETM is naturally described as a fiber bundle.
Analogies from the previous section can now be exploited. The wavenumber $k$ plays the role of the longitudinal coordinate $z$. The dispersion relation is analogous to the $\omega$-$z$ plot. Two specified Bloch modes $\mathbf{U}^{(1)}$ and $\mathbf{U}^{(2)}$ (e.g., corresponding to the longitudinal mode) correspond to the two curves of the $\omega$-$z$ plot. A physical classification scheme can be established using the following criteria: “Two waveguides belong to the same material class if they can be transformed into one another by continuous variation of either the geometric or the material properties without closing the gap between $\mathbf{U}^{(1)}$ and $\mathbf{U}^{(2)}$ in the dispersion relation.” This classification is also a topological classification of the fiber bundles. The topological invariant is the holonomy of parallel transport of a Bloch wave mode around the circular base manifold according to the Berry connection (Eq. (14) with $k$ replacing $z$). It is called the Zak phase, computed using Eq. (15) as
@@
The physical significance of the topological invariant manifests in the finite structure shown in Fig. 9g created by joining two waveguides with Zak phases equal to 00 and $\pi$, respectively. The resulting structure possesses a localized mode at the interface (Fig. 9h) at a frequency within the common band gap of the two halves of the composite waveguide (Fig. 9i). The mode is robust against small geometric or material perturbations that preserve the reflection symmetries of the structure as such perturbations do not change the topological class of the constituent waveguides.
@@ 7 Conclusions
This article reviewed the concept of geometric phase in the context of elastic waves and presented some applications to the design of dynamic and topological elastic materials. The geometric phase in quasi-1D waveguides was examined via practical examples, theoretical explanations, and design applications. The geometric phase was observed to depend solely on the path traced out in parameter space, and not on the rate of change of parameters. It can also be classified as either topological or nontopological. Only the former is robust against small perturbations to the design of the waveguide. The geometric phase was also illustrated using the concepts of fiber bundles and parallel transport, hence leading to its interpretation as the holonomy resulting from parallel transport. Similarly, the topological nature was distinguished based on the nature of the fiber bundle and its connection. Topological geometric phases require nontrivial fiber bundles resembling the edge of a Möbius strip, which are associated with full degeneracies in the parameter space describing the waveguide. Nontopological geometric phases require curved connections, which arise in rotating systems and in systems supporting polarized waves. The latter class of systems are described by the tangent bundle of a sphere. The interpretation of the geometric phase via arguments of differential geometry also enables its application to the design of ETMs. Indeed, ETMs can be described as fiber bundles classified according to topological invariants. Crucially, the abstract classification of fiber bundles indicates that the interface of ETMs with differing topological invariants can support localized edge modes. These edge modes are topologically protected against certain perturbations to the properties of the waveguide.
The concept of geometric phase provides a rigorous and solid foundation towards either the theoretical understanding of ETMs or the design of novel ETMs. Indeed, the geometric phase enables selective exploration of a vast design space in search of unique material properties such as those observed in topological material systems.
@@ Acknowledgements
M.K. thanks Prof. Ralph Kaufmann for useful discussions. M.K. acknowledges financial support from the Ross fellowship administered by the Purdue University Graduate School.

### v3

@@ 6 Topological classification of elastic waveguides
The first perspective establishes a classification scheme of the waveguides into two types, $\mathcal{W}_{O}$ and $\mathcal{W}_{E}$. Let the number of times the loop corresponding to a waveguide encircles the origin be denoted by $\nu$. Then, $\mathcal{W}_{O}$ ($\mathcal{W}_{E}$) is the class of the waveguide for which $\nu$ is odd (even). Waveguides in class $\mathcal{W}_{O}$ ($\mathcal{W}_{E}$) acquire (do not acquire) a geometric phase of $\pi$. For example, in Fig. 8a, the waveguide corresponding to loop $L_{1}$ ($L_{2}$) is in the class $\mathcal{W}_{O}$ ($\mathcal{W}_{E}$), which acquires (does not acquire) a geometric phase of $\pi$.
The properties of this classification scheme parallel the topological classification of metamaterials. Firstly, the classification scheme is robust against continuous geometric perturbations to a waveguide. Geometric perturbations to a given waveguide manifest as changes to the corresponding loop. If these perturbations are small enough, such as perturbing loop $L_{1}$ to $L_{1}^{\prime}$ in Fig. 8a, the waveguide class does not change. Secondly, waveguides in different classes are separated by gap closures in the $\omega$-$z$ plot. If a perturbation deforms a waveguide into a different class, there is an intermediate waveguide whose loop passes through the origin. For example, in Fig. 8a, perturbing the waveguide corresponding to loop $L_{1}$ into a waveguide corresponding to loop $L_{3}$ passes through an intermediate waveguide corresponding to loop $L_{2}$, for which the gap is closed in the $\omega$-$z$ plot.
Further, this classification scheme is also topological in nature. Recall from Sec. 5.1 that waveguides in class $\mathcal{W}_{O}$ ($\mathcal{W}_{E}$) give rise to a nontrivial (trivial) fiber bundle resembling the edge of a Möbius strip shown in Fig. 7e (edges of a cylinder, Fig. 7f). Thus, the classification scheme essentially distinguishes between these trivial and nontrivial fiber bundles. The topological invariant is the holonomy of the parallel transport along the base manifold, which is precisely the topological geometric phase (Figs. 7e,f).
@@ 6.2 Topological classification of elastic metamaterials
The above classification scheme is readily adapted to ETMs once the fiber bundle description is identified. 1D ETMs that are infinite and periodic along the $z$ direction are considered. Further, the unit cell of length $a$ is assumed to be symmetric about the $xz$ and $yz$ planes to prevent coupling between flexural, torsional, and longitudinal modes. Such metamaterials admit solutions of the form $\mathbf{u}(x,y,z,t)=\mathbf{U}(x,y,z;k)e^{\mathrm{i}(kz-\omega t)}$ to the NL equations, where $\mathbf{U}(x,y,z;k)$ is a periodic function of $z$ called the Bloch wave mode [64, 77]. The Bloch wave mode $\mathbf{U}(x,y,z;k)$ and angular frequency $\omega(k)$ are computed by solving a wavenumber dependent eigenvalue problem. These properties of the ETM are graphically described by plotting the dispersion relation, which shows the angular frequency as a function of the wavenumber, e.g., Figs. 9d-f. Due to the spatial periodicity of the metamaterial, the solutions are periodic in $k$ with periodicity $2\pi/a$, i.e., $\mathbf{U}(x,y,z;k)=\mathbf{U}(x,y,z;k+2\pi/a)$ and $\omega(k)=\omega(k+2\pi/a)$. In other words, the wavenumber is a circle $S^{1}$ represented as the interval $[-\pi/a,\pi/a]$ with the endpoints glued together. The wavenumber defines a circular base manifold. At each wavenumber $k$, the Bloch wave mode $\mathbf{U}(x,y,z;k)$ has a phase degree of freedom, as it can be arbitrarily multiplied by a phase factor $e^{\mathrm{i}\theta}$. This is described as a $U(1)$ fiber. Thus, a Bloch wave mode of an ETM is naturally described as a fiber bundle.
Analogies from the previous section can now be exploited. The wavenumber $k$ plays the role of the longitudinal coordinate $z$. The dispersion relation is analogous to the $\omega$-$z$ plot. Two specified Bloch modes $\mathbf{U}^{(1)}$ and $\mathbf{U}^{(2)}$ (e.g., corresponding to the longitudinal mode) correspond to the two curves of the $\omega$-$z$ plot. A physical classification scheme can be established using the following criterion: “Two waveguides belong to the same material class if they can be transformed into one another by continuous variation of either the geometric or the material properties without closing the gap between $\mathbf{U}^{(1)}$ and $\mathbf{U}^{(2)}$ in the dispersion relation”. This classification is also a topological classification of the fiber bundles. In this example, the topological invariant is the holonomy of parallel transport of a Bloch wave mode around the circular base manifold according to the Berry connection (Eq. (14) with $k$ replacing $z$). It is called the Zak phase, computed using Eq. (15) as
@@
The physical significance of the topological invariant manifests in the finite structure shown in Fig. 9g created by joining two waveguides with Zak phases equal to 00 and $\pi$, respectively. According to the bulk-boundary correspondence, the difference between the invariants of the constituent waveguides equals the number of localized edge modes at their interface. Here, the invariants differ by one unit because the Zak phase can only take values 00 and $\pi$. Thus, the resulting structure possesses a single edge mode at the interface (Fig. 9h) at a frequency within the common band gap of the two halves of the composite waveguide (Fig. 9i). The mode is robust against small geometric or material perturbations that preserve the reflection symmetries of the structure as such perturbations do not change the topological class of the constituent waveguides.
In general, for ETMs emulating the “periodic table” of topological metamaterials, the ETM maps to a fiber bundle with base manifold as the Brillouin zone (circle for 1D ETM, 2-torus for 2D ETM, 3-torus for 3D ETM) and fiber as the Bloch wave mode. The topological invariant is computed using standard formulae derived in the literature [26]. Note that the invariant is, in general, not a geometric phase. The invariant is determined by the dimension of the waveguide and its symmetries. The topological invariant relates to edge modes at the interface of topological distinct ETMs through the bulk-boundary correspondence.
ETMs based on more recent design principles such as higher-order topological metamaterials or topological defects also benefit from these concepts. Indeed, higher-order topological metamaterials are generalizations of the 1D Zak phase [72, 79] and phenomena at topological defects are extensions of the “periodic table” of topological metamaterials [73, 26].
@@ 7 Conclusions
This article reviewed the concept of geometric phase in the context of elastic waves and presented some applications to the design of dynamic and topological elastic materials. The geometric phase in quasi-1D waveguides was examined via practical examples, theoretical explanations, and design applications. The geometric phase was observed to depend solely on the path traced out in parameter space, and not on the rate of change of parameters. It can also be classified as either topological or nontopological. Only the former is robust against small perturbations to the design of the waveguide. The geometric phase was also discussed using the concepts of fiber bundles and parallel transport, hence leading to its interpretation as the holonomy resulting from parallel transport. Similarly, the topological nature was distinguished based on the nature of the fiber bundle and its connection. Topological geometric phases require nontrivial fiber bundles resembling the edge of a Möbius strip, which are associated with full degeneracies in the parameter space describing the waveguide. Nontopological geometric phases require curved connections, which arise in rotating systems and in systems supporting polarized waves. The latter class of systems are described by the tangent bundle of a sphere. The interpretation of the geometric phase via arguments of differential geometry also enables its application to the design of ETMs. Indeed, ETMs can be described as fiber bundles classified according to topological invariants. Crucially, the abstract classification of fiber bundles indicates that the interface of ETMs with differing topological invariants can support localized edge modes. These edge modes are topologically protected against certain perturbations to the properties of the waveguide.
The concept of geometric phase provides a rigorous and powerful tool to support either the theoretical understanding of ETMs or the design of novel ETMs. Indeed, the geometric phase complements other existing analysis tools to enable the selective exploration of the vast metamaterial design space in search of materials with unique properties.
@@ Acknowledgements
The authors gratefully acknowledge the partial financial support of the National Science Foundation under grant #2330957. M.K. thanks Prof. Ralph Kaufmann for useful discussions. M.K. acknowledges financial support from the Ross fellowship administered by the Purdue University Graduate School.